# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ 1.5 Notations
2 General Invariance Pursuit from Heterogeneous Environments
In this section, we formalize the problem of invariance pursuit using data from multiple environments. This formalization admits the canonical nonparametric invariance pursuit in Section 1.1 as a special case.
$\forall e\in\mathcal{E}\qquad\mathbb{E}\left[\left(Y^{(e)}-g^{\star}(X_{S^{\star}}^{(e)})\right)f(X^{(e)}_{S^{\star}})\right]=0\qquad\forall f\in[\Theta_{f}]_{S^{\star}},$
where $S^{\star}$ is the unknown set of true important variables. We refer to the above problem as invariance pursuit.
The problem of estimating $g^{\star}$ in (2) is a generalized version of the canonical nonparametric invariance pursuit with $g^{\star}=m^{\star}$ in (1.1) and $\Theta_{f}=\Theta_{g}=\Theta$. It depicts a general form and unifies several problems of interest in predecessors. For example, when $\Theta_{g}$ and $\Theta_{f}$ are all linear function classes, it reduces to the linear invariance pursuit problem, i.e., estimating $g^{\star}(x)=(\beta^{\star})^{\top}x=(\beta^{\star}_{S^{\star}})^{\top}x_{S^{\star}}$ with $\beta^{\star}\in\mathbb{R}^{d}$ satisfying $\mathrm{supp}(\beta^{\star})=S^{\star}$ in the multi-environment linear regression (Fan et al.,, 2023) with linear invariance structure
Another example is the augmented linear invariance pursuit where $\Theta_{g}$ is linear and $\Theta_{f}=\{f(x)=\sum_{j=1}^{d}\beta_{0,j}x_{j}+\beta_{1,j}\phi(x_{j})\}$ with some transform function $\phi:\mathbb{R}\to\mathbb{R}$. The augmented linear invariance structure that realizes (2) in this case is
$\mathbb{E}\left[\left(Y^{(e)}-(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}\right)X_{j}^{(e)}\right]=\mathbb{E}\left[\left(Y^{(e)}-(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}\right)\phi(X_{j}^{(e)})\right]=0\qquad\forall e\in\mathcal{E},j\in S^{\star}$
It coincides with the problem considered by Fan & Liao, (2014) when $|\mathcal{E}|=1$. Compared with the linear invariance pursuit that searches for the invariant best linear predictor across diverse environments while $\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]$ not necessary be linear, augmented linear invariance pursuit further leverages the structural knowledge that $\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]$ admits a near linear form, or the noise is at least uncorrelated with the transformed covariates in $S^{\star}$. Identification is possible in this case even when $|\mathcal{E}|=1$.
Similar to the discussion in Section 1.1, the main challenge here is the curse of endogeneity, that the population-level least squares minimizer in one environment $e\in\mathcal{E}$ diverges from $g^{\star}$, $g^{(e)}:=\mathop{\mathrm{argmin}}_{g\in\Theta_{g}}\mathbb{E}[|Y^{(e)}-g(X^{(e)})|^{2}]\neq g^{\star}$, the discrepancy can be arbitrary large; for a formal declaration, see Proposition 2 in Fan et al., (2023). To address this issue, we will harness the insight that the distributions of $(X,Y)$ across diverse environments capture the invariance structure (2). The central idea of this paper is to exploit both the heterogeneity among different environments, i.e., the shifts in population distributions $\mu^{(e)}$, in conjunction with the above invariance structure (2) to pinpoint the invariant regression function $g^{\star}$.
It should be noted that both $g^{\star}$ and $S^{\star}$ are determined by $(\Theta_{g},\Theta_{f})$ and $\mathcal{E}$ through the structure (2). It is required that $\partial\Theta_{g}=\{g-g^{\prime}:g,g^{\prime}\in\Theta_{g}\}\subseteq\Theta_{f}$. In the case of $\Theta_{f}=\partial\Theta_{g}$, one uses only heterogeneity among different environments, or the “invariance principle”, to identify the invariant regression function $g^{\star}$. Heterogeneous environments are essential in this case. By choosing substantially large $\Theta_{f}\supsetneq\partial\Theta_{g}$, one further injects the strong prior structural assumption that the invariant regression function lies in the class $\Theta_{g}$ rather than $\Theta_{f}\setminus\Theta_{g}$. In this case, one leverages both heterogeneity among environments, i.e., the “invariance principle”, and the mentioned prior structure knowledge, i.e., the “asymmetry principle”, to jointly identify $g^{\star}$. Only one environment may be enough for identifying $g^{\star}$ in some cases.
3 Focused Adversarial Invariance Regularized Estimator
In this section, we describe our proposed method of estimation: the Focused Adversarial Invariance Regularized Estimator. The estimation involves an optimization objective function, a linear combination of a user-designated risk function and our proposed focused adversarial invariance regularizer that exploits (2). The regularizer’s optimization involves a minimax/adversarial estimation among user-specified predictor function class $\mathcal{G}$ and discriminator function class $\mathcal{F}$ that approximate $\Theta_{g}$ and $\Theta_{f}$ respectively in Section 2. We first introduce the population-level regularizer to illustrate the underlying intuitions, which enforce the predictor to achieve fair optimal predictions, given the variables used, across diverse populations. The introduction of the complete framework follows this.
3.1 Imposing Invariance via Focused Adversarial Invariance Regularizer
We illustrate the idea of our regularizer for nonparametric invariance pursuit where $\Theta_{g}=\Theta_{f}=\Theta$. According to the discussion in Section 2, we hope to impose invariance structure akin to (1.1) on the predictor function $g$ we search in the class $\mathcal{G}$. That is, given the underlying ground truth $g$, we expect that
$\forall e\in\mathcal{E}\qquad\mathbb{E}[Y^{(e)}-g(X^{(e)})|X^{(e)}_{S_{g}}]=0.$ (3.1)
The above constraint is called conditional moment restriction, and is equivalent to
$\forall e\in\mathcal{E}\qquad\mathbb{E}\big{[}\{Y^{(e)}-g(X^{(e)})\}f(X^{(e)})\big{]}=0~{}~{}~{}~{}\forall\mbox{ testing function }f\in\Theta^{(e)}_{S_{g}}.$
Based on the above observation, we propose the following focused adversarial invariance regularizer. Let $\mathcal{F}$ be a function class that $\mathcal{F}\supseteq\mathcal{G}$. At the population level, it can be expressed as a minimax optimization objective function as
$\inf_{g\in\mathcal{G}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\underbrace{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[\big{\{}Y^{(e)}-g(X^{(e)})\big{\}}f^{(e)}(X^{(e)})-\frac{1}{2}\big{\{}f^{(e)}(X^{(e)})\big{\}}^{2}\right]}_{\mathsf{J}(g,f^{\mathcal{E}})}.$ (3.2)
The term “adversarial invariance” suggests that the candidate prediction function $g$ is penalized in a way such that the residual $Y^{(e)}-g(X^{(e)})$ defined by $g$ is uncorrelated with any function in $\mathcal{F}_{S_{g}}$. The term “focused” underscores that this penalty is applied in the direction of the selected variables of $g$. In other words, the variables that the test function $f^{(e)}$ works on are restricted within the variables used by $g$. The last $-\frac{1}{2}f^{2}$ term is incorporated to manage the scale automatically.
The above minimax optimization can be interpreted as a min-max game between the predictor $g\in\mathcal{G}$ and $|\mathcal{E}|$ discriminators $f^{\mathcal{E}}=(f^{(e)})_{e\in\mathcal{E}}\in\{\mathcal{F}\}^{|\mathcal{E}|}$. Intuitively, the discriminators from different environments act like police and attempt to discern whether the predictor satisfies the conditional moment restrictions within each environment $e\in\mathcal{E}$. On the other hand, the predictor, akin to a team of counterfeiters, strives to fool the discriminators across all the environments simultaneously.
Let us formally realize the above intuitions by showing that $\breve{g}$, the predictor in the solution of the above minimax objective with function classes $\mathcal{G}=\mathcal{F}=\Theta$, satisfies the conditional moment restriction (3.1). To this end, for a fixed $g\in\Theta$, by taking supremum over all the $f^{\mathcal{E}}\in\{\Theta_{S_{g}}\}^{\mathcal{E}}$, we obtain
$\mathsf{J}(g,\star):=\sup_{f^{\mathcal{E}}\in\{\Theta_{S_{g}}\}^{|\mathcal{E}|}}\mathsf{J}(g,f^{\mathcal{E}})$ $=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[\mathbb{E}\left[\sup_{u\in\mathbb{R}}\left\{\left(\mathbb{E}[Y^{(e)}|X_{S_{g}}^{(e)}]-g(X_{S_{g}}^{(e)})\right)u-\frac{1}{2}u^{2}\right\}\Big{|}X_{S_{g}}^{(e)}\right]\right]$
$=\frac{1}{2|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[|\mathbb{E}[Y^{(e)}|X_{S_{g}}^{(e)}]-g(X_{S_{g}}^{(e)})|^{2}\right].$
Given that $\mathsf{J}(g,\star)$ is non-negative, we can assert that the global minimal value of the objective $\mathsf{J}(g,\star)$ is 00, attained by the invariant regression function $m^{\star}$ in (1.1). Consequently, we can conclude that
$\breve{g}\in\mathop{\mathrm{argmin}}_{g\in\Theta}\mathsf{J}(g,\star)\qquad\Longleftrightarrow\qquad\breve{g}(X_{S_{g}}^{(e)})=\mathbb{E}[Y^{(e)}|X_{S_{\breve{g}}}^{(e)}=x_{S_{\breve{g}}}]~{}~{}\forall e\in\mathcal{E}.$ (3.3)
At the same time, the null $m^{\emptyset}(x)=0$ can also minimize $\mathsf{J}(g,\star)$ provided $Y$ is zero-mean across $\mathcal{E}$. It is easy to see if $S^{\star}$ can be divided into $K$ groups of mutually independent variables, then there are $2^{K}$ distinct global minima of $\mathsf{J}(g,\star)$. Here is where the risk loss $\ell(\cdot,\cdot)$ comes into play: it will select the most predictive variable set that preserves the invariance structure.
In practice, it is impossible to run the optimization program over large $\Theta_{g},\Theta_{f}$ given finitely many data, we will choose $\mathcal{G}\subseteq\Theta_{g}$ and $\mathcal{F}\subseteq\Theta_{f}$ that approximate $\Theta_{g},\Theta_{f}$, respective well. For example, when $\Theta_{g}=\Theta_{f}=\Theta$, we may let $\mathcal{G}$ and $\mathcal{F}$ be neural network classes with diverging $NL$.
3.2 An Unified Estimation Framework: FAIR Estimator
$\frac{\partial\ell(y,v)}{\partial v}=(v-y)\psi(v)\qquad\text{and}\qquad\frac{\partial^{2}\ell(y,v)}{\partial v^{2}}>0.$ (3.4)
The constraints in (3.4) ensure that the conditional expectation aligns with the unique global minima and can be satisfied by various risk losses. Two leading examples are the least square loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ with $\psi(v)=1$ for regression, and the cross-entropy loss $\ell(y,v)=-\log(1-v)-y\log\{v/(1-v)\}$ with $\psi(v)=1/\{v(1-v)\}$ for classification.
where $\mathcal{G}\subseteq\Theta_{g}$ and $\mathcal{F}\subseteq\Theta_{f}$ are function classes that approximates $\Theta_{g}$ and $\Theta_{f}$, respectively. Here $\widehat{\mathsf{R}}(g)$, the first part of the objective $\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})$, is the pooled empirical risk loss. This is defined as the pooled sample mean of the user-specified loss across all the environments $\mathcal{E}$:
$\widehat{\mathsf{R}}(g)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\widehat{\mathbb{E}}\left[\ell(Y^{(e)},g(X^{(e)}))\right]=\frac{1}{(|\mathcal{E}|\cdot n)}\sum_{e\in\mathcal{E}}\sum_{i=1}^{n}\ell(Y^{(e)}_{i},g(X^{(e)}_{i})).$ (3.6)
Integrating this loss prevents the estimator from collapsing to conservative solutions since both 00 and $m^{\star}$ attain the global minimum of the population-level regularizer. $\gamma$ is the hyper-parameter to be determined, and $\widehat{\mathsf{J}}(g,f^{\mathcal{E}})$ is the empirical counterpart of the regularizer in (3.2), that
$\widehat{\mathsf{J}}(g,f^{\mathcal{E}})=\frac{1}{(|\mathcal{E}|\cdot n)}\sum_{e\in\mathcal{E}}\sum_{i=1}^{n}\left[\big{\{}Y^{(e)}_{i}-g(X_{i}^{(e)})\big{\}}f^{(e)}(X^{(e)}_{i})-\frac{1}{2}\big{\{}f^{(e)}(X^{(e)}_{i})\big{\}}^{2}\right].$ (3.7)
$\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{\widehat{g}}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(\widehat{g},f^{\mathcal{E}})-(\gamma+1)\delta_{\mathtt{opt}}^{2}\leq\widehat{\mathsf{Q}}_{\gamma}(\widehat{g},\widehat{f}^{\mathcal{E}})\leq\inf_{g\in\mathcal{G}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})+(1+\gamma)\delta_{\mathtt{opt}}^{2}$ (3.8)
with some optimization error $\delta_{\mathtt{opt}}^{2}>0$, here $\gamma$ in $(1+\gamma)$ the same as that in $\widehat{\mathsf{Q}}_{\gamma}$?
4 Nonparametric Invariance Pursuit
In this section, we apply our proposed framework to the nonparametric invariance pursuit, whose data-generating process can be summarized as the following two conditions 4.1 and 4.2. In this case, we use deep ReLU networks with growing depth and width to realize the predictor and discriminator function class and choose the $L_{2}$ loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$, resulting in a Focused Adversarial Invariance Regularized Neural Network (FAIR-NN) least squares estimator.
From a methodological perspective, we use such a canonical estimator to illustrate the main idea and key features of our proposed framework. Firstly, one can seamlessly integrate black-box machine learning models (e.g. neural networks) into it and fully exploit these models’ sample efficiency and capability in being adaptive to low-dimension structures. Moreover, this framework allows to differentiate true important and spurious information from the covariates using only the data itself given its fully “nonparametric” nature. The above claims are supported by our theoretical analysis in Section 4.1 and Section 4.2, respectively.
From a theoretical perspective, Section 4.1 focus on the sample-efficiency of the pure statistical estimation problem nonparametric invariance pursuit: it shows that sample-efficiency (in both $n$ and $|\mathcal{E}|$) is viable under a minimal identification condition related to the heterogeneity of the environments. Meanwhile, Section 4.2 focus on the semantic meaning of $S^{\star}$ and $m^{\star}$ under the SCM with intervention setting. It shows that such an invariance structure (1.2) with some $S^{\star}$ always exists in this setting and the identification condition in Section 4.1 always holds for such a $S^{\star}$ under non-degenerate interventions. Such an identified $(m^{\star},S^{\star})$ has a simple and clear semantic meaning, and unveiling them from $\mathcal{E}$ has certain pragmatic utility. This provides a first yet comprehensive theoretical understanding of the nonparametric invariance pursuit.
@@ 4 Nonparametric Invariance Pursuit
To start with, we introduce some additional notations for theoretical analysis. Recall that $\mu^{(e)}$ is the joint distribution of $(X,Y)$ in environment $e$ and $\bar{\mu}=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}$ is the pooled joint distribution. We let $m^{(e,S)}\in\Theta_{S}^{(e)}$ be the conditional expectation of $Y$ given $X_{S}$ in environment $e$, i.e., $m^{(e,S)}(x)=\mathbb{E}[Y^{(e)}|X_{S}^{(e)}=x_{S}]$. Recall that $\nu_{x,S}$ is the marginal distribution of $X_{S}$ for $(X,Y)\sim\nu$. It is easy to see that $\mu^{(e)}_{x,S}$ is absolutely continuous with respect to $\bar{\mu}_{x,S}=[\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}]_{x,S}$ for any $S\subseteq[d]$, and we use $\rho^{(e)}_{S}$ to denote the Radon–Nikodym derivative of $\mu_{x,S}^{(e)}$ with respect to $\bar{\mu}_{x,S}$. We define $\bar{m}^{(S)}(x)=\sum_{e\in\mathcal{E}}\rho^{(e)}_{S}(x_{S})m^{(e,S)}(x)$, which can be interpreted as the population-level least-squares that regress $Y$ on $X_{S}$ using all the data.
Condition 4.1 (Data Generating Process).
Condition 4.2 (Invariance and Nondegenerate Covariate).
@@
There exists some set $S^{\star}$ and $m^{\star}:\mathbb{R}^{|S^{\star}|}\to\mathbb{R}$ such that $m^{(e,S^{\star})}(x)\equiv m^{\star}(x_{S^{\star}})$ for any $e\in\mathcal{E}$. Moreover, for any $S\subseteq[d]$ with $S^{\star}\setminus S\neq\emptyset$, $\inf_{m\in\Theta_{S}}\|m-m^{\star}\|_{2}^{2}\geq s_{\min}>0$.
@@
As stated in 4.1, we consider a more general setting where only the data within each environment are independent; data across different environments may be dependent. In 4.2, the “Invariance” condition is just a restatement of (1.1). The “Nondegenerate Covariate” condition rules out some degenerate cases, for example, $m^{\star}(x_{1})=x_{1}^{2}$ with $S^{\star}=\{1\}$ and $X_{2}=X_{1}^{4}$, or $m^{\star}(x_{1},x_{2})=x_{1}$ with $S^{\star}=\{1,2\}$, and is imposed for technical convenience. The target (invariant) regression function in nonparametric invariance pursuit is $m^{\star}$.
We also introduce the following two standard regularity conditions in nonparametric regression.
Condition 4.3 (Sub-Gaussian Response).
Condition 4.4 (Boundedness in Nonparametric Regression).
We consider realizing the predictor function class and discriminator function class as
$\mathcal{G}=\mathcal{H}_{\mathtt{nn}}(d,L,N,b_{m})\qquad\text{and}\qquad\mathcal{F}=\mathcal{H}_{\mathtt{nn}}(d,L+2,2N,2b_{m})$ (4.1)

### v2

@@ 1.6 Notations
2 FAIR Least Squares Estimator Using Neural Networks
In this section, we show that one can use the FAIR-NN least squares estimator, a realization of the FAIR estimator by setting $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ and specifying both $(\mathcal{G},\mathcal{F})$ as neural networks, to attain sample-efficient estimation in nonparametric causality pursuit.
The main messages of this section are two-fold. From a theoretical perspective, it shows that sample-efficient estimation (in both $n$ and $|\mathcal{E}|$) in the general nonparametric causality pursuit problem is viable under a minimal identification condition related to the heterogeneity of the environments. From a methodological perspective, it demonstrates one key feature of our proposed framework: one can seamlessly integrate black-box machine learning models (e.g. neural networks) into it and fully exploit these models’ sample efficiency and capability in being adaptive to low-dimension structures.
2.1 Setup
@@
We introduce some notations. Recall that $\mu^{(e)}$ is the joint distribution of $(X,Y)$ in environment $e$. Let $m^{(e,S)}(x):=\mathbb{E}[Y^{(e)}|X_{S}^{(e)}=x_{S}]$ be the conditional expectation of $Y$ given $X_{S}$ in environment $e$. Recall that $\nu_{x,S}$ is the marginal distribution of $X_{S}$ for $(X,Y)\sim\nu$. It is easy to see that $\mu^{(e)}_{x,S}$ is absolutely continuous with respect to $\bar{\mu}_{x,S}=[\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}]_{x,S}$ for any $S\subseteq[d]$ hence $\rho^{(e)}_{S}$, the Radon–Nikodym derivative of $\mu_{x,S}^{(e)}$ with respect to $\bar{\mu}_{x,S}$, is well defined. We define $\bar{m}^{(S)}(x)=\sum_{e\in\mathcal{E}}\rho^{(e)}_{S}(x_{S})m^{(e,S)}(x)$, which can be interpreted as the population-level least squares that regress $Y$ on $X_{S}$ using all the data in $\mathcal{E}$.
Condition 1 (Model and Regularity Conditions).
There exists some positive constants $(C_{0},s_{\min})$ such that the following conditions hold.
  • (a)
    Data Generating Process We collect data from $|\mathcal{E}|\in\mathbb{N}^{+}$ environments with $|\mathcal{E}|\leq n^{C_{0}}$. For each environment $e\in\mathcal{E}$, we observe $\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{i=1}^{n}\overset{i.i.d.}{\sim}\mu^{(e)}$.
  • (b)
@@
    Invariance Structure: There exists some set $S^{\star}$ and $m^{\star}:\mathbb{R}^{|S^{\star}|}\to\mathbb{R}$ such that $m^{(e,S^{\star})}(x)\equiv m^{\star}(x_{S^{\star}})$ for any $e\in\mathcal{E}$.
  • (c)
    Sub-Gaussian Response: For any $e\in\mathcal{E}$ and $t\geq 0$, $\mathbb{P}\left[|Y^{(e)}|\geq t\right]\leq C_{0}e^{-t^{2}/(2C_{0})}$.
  • (d)
    Boundedness: $X\in[-C_{0},C_{0}]^{d}$ $\bar{\mu}$-a.s. and $\|m^{(e,S)}\|_{\infty}\leq C_{0}$ for any $S\subseteq[d]$ and $e\in\mathcal{E}$.
  • (e)
    Nondegenerate Covariate: For any $S\subseteq[d]$ with $S^{\star}\setminus S\neq\emptyset$, $\inf_{m\in\Theta_{S}}\|m-m^{\star}\|_{2}^{2}\geq s_{\min}>0$.
@@
1 (a)–(b) is just a restatement of (1.1) together with i.i.d. data within each environment; data across different environments may be dependent. (c)–(d) are standard in nonparametric regression. (e) rules out some degenerate cases, for example, $m^{\star}(x_{1})=x_{1}^{2}$ with $S^{\star}=\{1\}$ and $X_{2}=X_{1}^{4}$, or $m^{\star}(x_{1},x_{2})=f(x_{1})$ with $S^{\star}=\{1,2\}$, and is imposed for technical convenience. The target (invariant) regression function in nonparametric causality pursuit is $m^{\star}$.
2.2 Proposed FAIR-NN Least Squares Estimator
Given all the data $\{\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{i=1}^{n}\}_{e\in\mathcal{E}}$ from heterogeneous environments, we consider using the following FAIR-NN least squares estimator to learn $m^{\star}$ in (1.1). Specifically, the FAIR-NN least squares estimator is the solution to the subsequent minimax optimization objective
$\widehat{g}\in\mathop{\mathrm{argmin}}_{g\in\mathcal{G}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\frac{1}{|\mathcal{E}|\cdot n}\sum_{e\in\mathcal{E},i\in[n]}\left\{Y^{(e)}_{i}-g(X^{(e)}_{i})\right\}^{2}+\gamma\widehat{\mathsf{J}}(g,f^{\mathcal{E}}).$ (2.1)
where the first part of the objective $\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})$ is the pooled least squares loss preventing the estimator from collapsing to conservative solutions, $\gamma$ is the hyper-parameter to be determined, and $\widehat{\mathsf{J}}(g,f^{\mathcal{E}})$ is the empirical counterpart of the focused adversarial invariance regularizer defined as
$\widehat{\mathsf{J}}(g,f^{\mathcal{E}})=\frac{1}{|\mathcal{E}|\cdot n}\sum_{e\in\mathcal{E},i\in[n]}\left[\big{\{}Y^{(e)}_{i}-g(X_{i}^{(e)})\big{\}}f^{(e)}(X^{(e)}_{i})-\frac{1}{2}\big{\{}f^{(e)}(X^{(e)}_{i})\big{\}}^{2}\right].$ (2.2)
The minimax program (2.1) is the empirical version of (1.2) via setting $\ell(y,v)=\frac{1}{2}(y-v)^{2}$. Here we specify the predictor function class $\mathcal{G}$ and testing (discriminator) function class $\mathcal{F}$ as
$\mathcal{G}=\mathcal{H}_{\mathtt{nn}}(d,L,N,B)\qquad\text{and}\qquad\mathcal{F}=\mathcal{H}_{\mathtt{nn}}(d,L+2,2N,2B)$ (2.3)